3. Find V. 8 ΚΩ ΕΛΛΗ 2 ΚΩ 4 ΚΩ 4 ΚΩ 9 V 2v +
Answer #1

2 M Ukl Vi 4kn V3 Vo 9V 2 short circuit V3 = av Due to virtuel and Va=vo. At NODE Eeln vi-q VI-V3 vina + vi- va + -o put V₂=2solve Ealo 5V 22 +252 6 Equation 3 multiply with ech x 5 5v/ +10v2 = 30 -V2 = SVI 22 20 + 8 11V2 8 -0.7272 V2 = Vo Answer Vo